Make New Entry,
Make Followup Entry
User name Armstrong
Log entry time 10:06:06 on November30,2009
Entry number 303679
keyword=Cryogen flow blip
The Hall is in controlled access, for Heidi, Todd, and Howard to
work on recovering the right HRS magnets.
The target has been happy, however there was a cryogen flow
fluctuation, presumably related to their work on the magnets.
The fluctuation is at -19 minutes on this strip chart.
For the record, the target is loop 1, 20 cm LD2.
Question - I have not adjusted down the JT valve while the
beam is off. The HP heater is on, at about 1030 W. Should the protocol
be to (slowly) close the JT when we know that beam will be off for an
extended time (i.e. for a long controlled access), in order to reduce
the cooling demands on ESR